A car travels 400 miles on level terrain in the same amount of time it travels 160 miles on mountainous terrain. If the rate of the car is 30 miles per hour less in the mountains than on level ground, find its rate in the mountains.
Found 2 solutions by nyc_function, mananth:
Answer by nyc_function(2741)

Here is your equation:
400/(x + 30) = 160/(x - 30)

Can you finish?

Answer by mananth(16949)

rate of car on level ground be x mph
on the mountains it will be x-30 mph
...
distance traveled on level ground = 400 miles
distance traveled in the mountains = 160 miles
...
t = d/r
..
time is sdame on both legs.
..
400/x=160/(x-30)
400(x-30)=160x
400x-12000=160x
+12000 +12000
400x = 160x+12000
-160x -160x
240x=12000
/240
240x/240 = 12000/240
x=50 mph on level ground
in the mountains it will go at 50-30 = 20 mph
...
m.ananth@hotmail.ca